A heat exchanger is to be designed for the following specifications:         Hot gas temperature, 1145...

  1. A heat exchanger is to be designed for the following specifications:

        Hot gas temperature, 1145 oC

        Cold gas temperature, 45 oC

        Unit surface conductance on the hot side, 230 W/m2-K

        Unit surface conductance on the cold side, 290 W/m2-K

        Thermal conductivity of the metal wall, 115 W/m-K

Find the maximum thickness of the metal wall between the hot gas and cold gas so that the    maximum temperature of the wall does not exceed 545 oC.
Answer #1

T Tz hcold = 2.90 W/m²K hhot = 230 w/m kl That = 1145°c Tcold - 45°C - Th=545c K=115 W/m-k For the convection heat transfer